OpenStudy (newwar):

What is the reason for each step in the solution of the equation? 12−5x+6x=412−5x+6x=4 Drag and drop the reasons into the boxes to correctly complete the table. 12−5x+6x=412−5x+6x=4 12+x=412+x=4 x=−8x=−8 GivenDistributive PropertyCombine Like TermsSubtraction Property of EqualityAssociative Property

OpenStudy (sshayer):

12-5x+6x=4 subtract 12 from both sides 12-5x+6x-12=4-12 -5x+6x=-8 x=-8